The odor of spoiled butter is due in part to butanoic acid, which results from the chemical breakdown of butter fat. A 0.100 M solution of butanoic acid is 1.23% ionized. Calculate the equilibrium constant for this weak monoprotic acid.
Answer #1

Ionization reaction :

CH3CH2CH2COOH ​​+ H2O <==> H3O+   + CH3CH2CH2COO-

Equilibrium constant(kc) is written as :

kc=[H3O+ ][CH3CH2CH2COO-]/[CH3CH2CH2COOH][H2O]----(1)

Acid is 1.23% ionized so there is 98.77 unionized acid.

[H+ ] = [CH3CH2CH2COO-] = 0.0123(0.100M)=0.00123 M

[CH3CH2CH2COOH]=0.9877(0.100M)=0.09877 M

[H2O]= 1M

Putting all these values in equation (1)

kc= [H+ ][CH3CH2CH2COO-]/[CH3CH2CH2COOH][H2O]

kc =(0.00123M)(0.00123M)/(0.09877M)(1M)

kc = 0.0000015129/0.09877

kc = (1.5129 x 10-6)/0.09877

kc = 0.153 x 10 -4

Equilibrium constant of weak monoprotic acid is 0.153 x 10 -4